A class is given a list of 20 study problems from which 10 will be chosen as part of an upcoming exam. A given student knows how to solve 15 of the problems. Find the probablility that the student will be able to answer (a) all 10 of the questions on the exam, (b) exactly 8 questions on the exam, and (c) at least 9 questions on the exam.
Something I for (a) was
probability that he will know the answer to one question=15/20=.75
(.75)^10=.056
the answers were
(a).016 or 21/1292 (b) .348 or 225/646 (c) .152 or 49/323

Answer by stanbon(75887)

A class is given a list of 20 study problems from which 10 will be chosen as part of an upcoming exam. A given student knows how to solve 15 of the problems. Find the probablility that the student will be able to answer
------
(a) all 10 of the questions on the exam,
# of ways to succeed: 15C10 = 3003
# of possible outcomes: 20C10 = 184756
---
P(10 correct) = 3003/184756 = 0.016
=========================================
(b) exactly 8 questions on the exam, and
# of ways to succeed: 15C8*5C2 = 6435*10 = 64350
# of possible outcomes: 184756
P(8 correct) = 64350/184756 = 0.0348
----------------------------------------
(c) at least 9 questions on the exam.
the answers were
[15C9*5C1 + 15C10]/20C10 = 0.152
==================================
Cheers,
Stan H.